The power meter used in the measurements of elog entries 2822, 3698 and 3709 was the Ophir PD300-3W. This power head has several damaged patches and a slight movement of the laser spot changes the reading considerably. To verify I checked the power out with another power meter (the Vector S310) and found that there is no significant variation of the power output with the temperature of the laser. And the power at 2.1A of diode current is 2W with 10% fluctuation arising from slight repositioning of the laser head. There are regions of the Ophir PD300 which show the laser power to be about 1.9W.